OpenStudy (anonymous):

Verify the trigonometric equation by making the left side match the right:-tan^2x+sec^2x=1

OpenStudy (loser66):

break tan = sin/cos sec = 1/ cos make them have the same denominator and you can get the right hand side Show your brain,

OpenStudy (anonymous):

I am still confused. How do I make the left and right side match?

OpenStudy (loser66):

i gave you the way, just apply.

OpenStudy (anonymous):

Ok, so now I have -(sin x/cos x)^2+(1/cos x)^2=1 correct?

OpenStudy (loser66):

you don't have 1 from the right hand side, you are trying to get it. Therefore, you cannot write it down, Take it off

OpenStudy (anonymous):

I was just carrying the right hand side down. I know it does not equal it yet. I believe I need to use the Pythagorean Identity sin^2x+cos^2x=1 now, but I am not sure what to do with it.

OpenStudy (loser66):